Suppose Coulomb measures a force of 2.0×10−5N between the two charged spheres when they are separated by 5.0cm . By turning the dial at the top of the torsion balance, he approaches the spheres so that they are separated by 3.0cm . What force does he measure now?
Answer #1

From coulombs law

F= k*q1*q2/(r^2)

when they are separated by 5.0 cm

Finitial =k*q1*q2/(ri^2)

when they are separated by 3.0 cm

Ffinal =k*q1*q2/(rf^2)

Ffinal /(Finitial) =(ri^2)/(rf^2)

Ffinal = Finitial *(ri/rf)^2

= 2*10^-5 N *(5 cm/3 cm)^2

=5.6*10^-5 N